OpenStudy (anonymous):

how would i write -9.38 X 10^5/2 X 10^-4 in scientific notation

OpenStudy (mertsj):

\[\frac{-9.38 \times 10^5}{2\times10^{-4}}= -4.69\times10^9\]

OpenStudy (anonymous):

thanks man

OpenStudy (mertsj):

yw